OpenStudy (anonymous):

Find the inverse of the function. f(x) = x3 - 8

OpenStudy (redohawk):

choices?

OpenStudy (nirmalnema):

let y=x^3-8 x^3=y+8 \[x=\sqrt[3]{y+8}\] now simply interchange x and y \[y=\sqrt[3]{x+8}\] is its inverse

OpenStudy (nirmalnema):

@Ale01jandra

OpenStudy (nirmalnema):

do u get that

OpenStudy (anonymous):

yes I do @nirmalnema thank you !

OpenStudy (nirmalnema):

ok ...!! your welcome.. !! :)